OpenStudy (anonymous):

Use Property 1 for radicals to simplify the following radical expression as much as possible. Assume the variable represents a positive number.

OpenStudy (anonymous):

\[\sqrt{50x^2}\]

hartnn (hartnn):

\(= \sqrt {50} \sqrt {x^2}\)

hartnn (hartnn):

\(= \sqrt 2\sqrt {25} \sqrt {x^2}\)

OpenStudy (anonymous):

the middle is 5 & the last one is x

hartnn (hartnn):

correct! \(= 5\sqrt2x\)

hartnn (hartnn):

thats it!

OpenStudy (anonymous):

Thanks!

hartnn (hartnn):

welcome ^_^